which of the following does not represent the sides of a right triangle?3,4,5 4,5,6 5,12,13 or 6,8,10

Answer by Edwin McCravy(20054)

We check to see if the Pythagorean theorem is satisfied.
The LONGEST side will be the HYPOTENUSE, c.  It doesn't
matter which of the other two sides is a and which is b.

3,4,5 

c=5, a=3, b=4 

Substitute in

c² = a²+b²
5² = 3²+4²
25 = 9+16
25 = 25   That's true so 3,4,5 form a right triangle.

--------------------------------------------------

4,5,6 

c=6, a=4, b=5 

Substitute in

c² = a²+b²
6² = 4²+5²
36 = 16+25
36 = 41   That's false so 4,5,6 does not form a right triangle.

--------------------------------------------------

5,12,13 

c=13, a=5, b=12 

Substitute in

c² = a²+b²
13² = 5²+12²
169 = 25+144
169 = 169   That's true so 5,12,13 form a right triangle.

--------------------------------------------------

6,8,10

c=10, a=6, b=8 

Substitute in

c² = a²+b²
10² = 6²+8²
100 = 36+64
100 = 100   That's true so 6,8,10 form a right triangle.

Edwin